OpenStudy (anonymous):

Which of the following options results in a graph that shows exponential growth? A.f(x) = 0.4(3)x B.f(x) = 3(0.5)x C.f(x) = 0.8(0.9)x D.f(x) = 0.9(5)−x

OpenStudy (anonymous):

@ali2x2 can u help me plz

OpenStudy (anonymous):

Increasing the value of x, which of these functions will get bigger

OpenStudy (anonymous):

?

OpenStudy (anonymous):

okay whats next

OpenStudy (anonymous):

let's check b., \[f(x)=3(0.5)^x\], that is the same as: \[f(x)=3(\frac{1}{2})^x = \frac{3}{2^x}\] so if we increase the value of x, f(x) will became each time smaller till it reaches near zero as x goes to infinity

OpenStudy (anonymous):

The same occurs for c.

OpenStudy (anonymous):

so B would be wrong

OpenStudy (anonymous):

With d. we have: \[f(x)=0.5(9)^{-x} = 0.5(\frac{1}{9})^x = \frac{0.5}{9^x}\]

OpenStudy (anonymous):

same thing happens here

OpenStudy (anonymous):

so, the only one that will increase with x will be...

OpenStudy (anonymous):

A

OpenStudy (anonymous):

;)

OpenStudy (anonymous):

thx u are a big help

OpenStudy (anonymous):

No problem ^^